# CTP-SCU/2025004

### v1

@@ Abstract
We present an operator-algebraic definition for timelike entanglement entropy in QFT. This rigorously defined timelike entanglement entropy is real-valued due to the timelike tube theorem. A simple proof of the $F$-theorem is provided by the strong subadditivity inequality of this timelike entanglement entropy.
@@ Introduction
A simple proof of $F$-theorem
We now turn to the dynamics of timelike entanglement entropy under the renormalization group (RG) flow. For $(2+1)$-dimensional QFTs, the $F$-theorem [19, 20, 21, 22] states that:
  1. 1.
    There exists a function $\mathcal{F}$ that equals the sphere free energy $F$ associated with the CFT at each fixed point of RG flows.
  2. 2.
    $\mathcal{F}$ monotonically decreases along RG flows.
An entropic $\mathcal{F}$-function [23] was introduced as
$\mathcal{F}=(R\partial_{R}-1)S(V),$ (16)
where $S(V)$ is the entanglement entropy of a disk $V$ with radius $R$. The monotonicity of $\mathcal{F}$-function can be proved by the strong subadditivity inequality of entanglement entropy $S(V)$ [24], although technical details of the proof are somewhat intricate.
We are interested in the $\mathcal{F}$-function of timelike entanglement entropy $S(\mathcal{T})$, or $S(T)$ for clarity. At a fixed point of the RG flow, using $S(V)|_{\mathrm{CFT}_{3}}=\alpha\pi\frac{2R}{\epsilon}-F$, from equation (14), we have
$S(T)|_{\text{CFT}_{3}}=\alpha\frac{\pi T}{\epsilon}-F,$ (17)
where $\alpha$ is the model-dependent constant. Set an entropic function
$\mathcal{F}=(T\partial_{T}-1)S(T),\quad\mathcal{F}|_{\text{CFT}_{3}}=F.$ (18)
The key point is that $S(T)$ depends only on the timelike interval, which enables us to simplify the strong subadditivity inequality of timelike entanglement entropy.
Figure 3: One timelike interval is divided into three parts $\mathcal{X}$, $\mathcal{Y}$ and $\mathcal{Z}$, where $\mathcal{Y}$ and $\mathcal{X}\mathcal{Y}\mathcal{Z}$ correspond to two centric spacelike disks.
Suppose we divide a timelike interval into three parts $\mathcal{X}$, $\mathcal{Y}$ and $\mathcal{Z}$, as depicted in Figure 3. Specifically, we choose three intervals as
$\mathcal{X}$ $=$ $\left\{\frac{T}{2}<t<\frac{T}{2}+\Delta T,\vec{x}=\vec{x}_{0}\right\},$ (19)
$\mathcal{Y}$ $=$ $\left\{-\frac{T}{2}<t<\frac{T}{2},\vec{x}=\vec{x}_{0}\right\},$ (20)
$\mathcal{Z}$ $=$ $\left\{-\frac{T}{2}-\Delta T<t<-\frac{T}{2},\vec{x}=\vec{x}_{0}\right\}.$ (21)
The strong subadditivity inequality requires
$S(\mathcal{XY})+S(\mathcal{YZ})\geq S(\mathcal{XYZ})+S(\mathcal{Y}),$ (22)
which leads to
$S(T+2\Delta T)+S(T)-2S(T+\Delta T)\leq 0.$ (23)
Taking the limit $\Delta T\rightarrow 0$, we obtain
$S^{\prime\prime}(T)\leq 0,$ (24)
which indicates the monotonicity of $\mathcal{F}$-function:
$\mathcal{F}^{\prime}=TS^{\prime\prime}(T)\leq 0.$ (25)
Therefore, by applying the strong subadditivity of timelike entanglement entropy, we provide a simple proof of $F$-theorem.
Figure 4: In two dimensions, spacelike intervals can be related to timelike intervals by a $\frac{\pi}{2}$-rotation.
In $(1+1)$-dimensional QFTs, the entropic analogue to Zamolodchikov’s $C$-theorem [25] can be also proved by the strong subadditivity inequality for entanglement entropy of spacelike intervals [26, 27]. We can define the $C$-function for timelike entanglement entropy $S(T)$ as
$C=3TS^{\prime}(T),$ (26)
and prove the monotonicity of entropic $C$-function with the same fashion in [27], because of a perfect symmetry between space and time, as shown in Figure 4.
@@
In this letter, using known results in algebraic QFT, we proposed a rigorous definition for timelike entanglement entropy in QFT of general dimensions. The key insight is that timelike entanglement entropy depends only on the timelike interval and is real-valued. As an application, we demonstrated that the strong subadditivity of timelike entanglement entropy provides a simple proof of $F$-theorem.
@@
In this work, the simplicity of the proof for $F$-theorem in QFT3 is attributed to the dimensionality of one in the temporal direction. This nice property holds for all QFTd. To prove the similar theorems in $d>3$, the only challenge is constructing appropriate entropic functions.
@@ Acknowledgments.
This work is supported in part by NSFC (Grant No. 12105191 and 12275184).

### v2

@@ Abstract
We present an operator-algebraic definition for timelike entanglement entropy in QFT. This rigorously defined timelike entanglement entropy is real-valued due to the timelike tube theorem.
@@ Introduction
In this letter, using known results in algebraic QFT, we proposed a rigorous definition for timelike entanglement entropy in QFT of general dimensions. The key insight is that timelike entanglement entropy depends only on the timelike interval and is real-valued.
@@ Acknowledgments.
We deeply appreciate Edward Witten for pointing out that the strong subadditivity fails by the non-commutativity of the algebras in the timelike direction. This work is supported in part by NSFC (Grant No. 12105191 and 12275184).